OpenStudy (anonymous):

Choose the number that is 1/4 of 1/2 of 1/5 of 200:

OpenStudy (rc94):

this is essentially just multiplying those numbers together, so you get 5

OpenStudy (anonymous):

of means multiply therefore, 1/4 * 1/5 * 1/2 = 1/40 1/40 multiplied by 200 = 5